RC meeting and Run Plan, Monday, March 24th

From SBS wiki
Jump to: navigation, search

Back to the calendar

RC Meeting Information

Current RC: Xinzhan Bai

RC Daily Meetings are at 3:30 (pm) via Zoom and in person in the counting house meeting room (2nd floor)

Useful Links

Accelerator

  • Accelerator main focus this week is still RF recovery
  • Facility working on building 91 and 95, installed pumps on 91, pumps working. Sealing contractor getting badge today, work to be completed by Wednesday
  • trouble shooting 1L09, trying to bring it up online
  • Beam expected on Wednesday for Hall C, Hall A follow a delayed beam delivery. Hall C shifts canceled for Wednesday OWL, day shift on standby

Resolved Issues

  • ECal Heating test - ( succeeded ) ECAL Test Results
  • Alarm handler updated, CDET HV added to alarm handler ( Done. see this link)
  • SBS Counter Platform cleared ( ready for DC power test on Wednesday)
  • Safety Documents done.

Current work in the Hall

  • DC power HC of the Hall arc dipole and BSY - Wednesday afternoon (Note: clean up all small items away out of the Hall by Wednesday )
  • Gate valve work near downstream of target - Monday
  • Survey on Target - Tuesday
  • Lead wall installation
  • Scintillator pedals coming down Today
  • Beam authorization in progress

TO Training

  • Hall C target cooled down, reach out to target group see if you need TO practical training

Shifts

  • Shifts canceled through March 30 following beam schedule

Beam requirements

  • start off at 3 pass beam
  • intrinsic beam size: 500 um
  • raster size: 2 mm X 2 mm; using carbon hole target with 2mm hole to checkout raster size
  • For later
    • production beam current : 50 uA

SBS Magnet

  • Note: SBS added more coils
  • First the first 3 pass, SBS magnet will be at 41.5% of maximum
  • corrector magnet setting for each pass passed to Yves Roblin

Status of Subsystem

    • ECal
    • HCal
    • GEMS
      • Gas system
    • CDET
    • DAQ
    • Analysis Software, online monitor, 50k replay
    • ...

Current Issues

  • Experts to update wiki for shift takers (DAQ/GEM/CDET/HCAL/ECAL/50k replay)
  • Update zoom software on the computer in conference room 200C